The number of goals Alice scores in a soccer game is a Poisson(2.2) random variable. Her coach awards her 2k candies if she scores k goals in a game. For example, if she scores 1 goal she gets 2 candies, if she scores 2 goals she gets 4 candies, if she scores 3 goals she gets 8 candies, and so on. If Alice does not score any goals, she does not get any candies. What is the expected number of candies Alice gets in a game?
1
Expert's answer
2020-11-03T17:15:26-0500

P(k goals in match) = (2.2k*exp[-2.2])/k!

expected number of candies Alice gets in a game = "\displaystyle\sum_{k=1}^\infty" [(2.2k*e-2.2)/k!*2k]=

"\displaystyle\sum_{k=1}^\infty" [(2.2k*e-2.2)/k!*2k] = e-2.2 "\displaystyle\sum_{k=1}^\infty" [(4.4k)/k!] = e-2.2 ( "\displaystyle\sum_{k=0}^\infty" [(4.4k)/k!] - 4.40/0!) = e-2.2 (e4.4 - 1) = 8.91